Grade 12th passIIT JEE Entrance Exam

A bulb is made using 2 filaments . A switch selects whether the filament are used individually or in parallel . When used with a 15V battery , the bulb can be operated at 5W, 10W, or 15W. What should be the resistances of the filaments ? I got the answers for 5W nd 10W but the answer for 15W is given to be not possible...why?

Abbas
11 Years agoGrade 12th pass

1 Answer

Arun
6 Years ago
Let R1 and R2 be the resistances of the filaments in the bulb individually and R3 be the equivalent resistance of the two resistances which is selected by the switch.
Given :
V = 15 V
Case 1 :
When the switch selects first resistance R1, then power of the bulb, P1 = 5 W
Using formula :
P = V2R
R1 = V2P1we get R1 = (15)25 = 49 Ω
Case 2 :
 
When the switch selects second resistance R2, then power of the bulb, P2 = 10 W
​using above formula, we get R2 = 22.5 Ω
 
Case 3 :
When the switch selects the filaments of resistances R1 and R2 in parallel connection , then the equivalent resistance of the both resistances is R3 and power of the bulb P3 = 15 W
​R3 = V2/P3 = 225/15 = 15 ​ Ω